### v1

# Infinitely many counterexamples to a conjecture of Lovász

Abstract
Motivated by the well-known conjecture of Ryser which relates maximum matchings to minimum vertex covers in $r$-partite $r$-uniform hypergraphs, Lovász formulated a stronger conjecture. It states that one can always reduce the matching number by removing $r-1$ vertices. This conjecture was very recently disproven for $r=3$ by Clow, Haxell, and Mohar using the line graph of a $3$-regular graph of order $102$. Building on this, we describe a simple infinite family of counterexamples based on generalized Petersen graphs for the case $r=3$ and give specific counterexamples for $r=4$.
1 Introduction
Let $H$ be an $r$-uniform hypergraph. The matching number $\nu(H)$ is the maximum number of pairwise disjoint edges and the vertex cover number $\tau(H)$ is the minimum cardinality of a set of vertices intersecting every edge of $H$. Furthermore, we say that $H$ is $r$-partite if there exists a partition of the vertex set of $H$, $V_{H}=V_{1}\cup\cdots\cup V_{r}$, such that $|e\cap V_{i}|=1$ for every $e\in E_{H}$ and $1\leq i\leq r$. It follows immediately from the above definitions that
$\nu(H)\leq\tau(H)\leq r\,\nu(H).$
In general both inequalities are easily shown to be tight. However, for the particular case of $r$-partite $H$, Ryser conjectured, first explicitly stated in Henderson’s 1971 dissertation [Hen71], that
$\tau(H)\leq(r-1)\,\nu(H).$
For $r=2$ this is true do to Kőnig’s theorem. Beyond this, the conjecture is only known to hold for the case $r=3$ [Aha2001] as well as special families of hypergraphs [FraHerMcKWan17, BisDasMorSza21].
A potential proof strategy for Ryser’s conjecture consists of constructing the vertex cover by iteratively removing $r-1$ vertices in such a way that it decreases the matching number. Lovász, in his 1975 dissertaton [Lov75] and later popularized through a 1988 survey of Füredi [furedi1988matchings], conjectured that this should in fact be possible.
Conjecture 1.
(Lovász) If $H$ is an $r$-partite $r$-uniform hypergraph and contains at least one edge, then there exists a set of $r-1$ vertices whose deletion reduces its matching number.
[…]
Figure 1: The $3$-regular generalized Petersen graph $\operatorname{GP}(11,2)$ on $22$ vertices whose line hypergraph is a counterexample for Lovász’ conjecture for $r=3$. It generalizes to the infinite family of graphs $\operatorname{GP}(5k+11,2)$ whose line hypergraphs are counterexamples to Lovász’ conjecture for $r=3$ and any $k\geq 0$.
The case $r=2$ also follows from Kőnig’s theorem, since removing any vertex from a minimum vertex cover creates a graph with vertex cover number, and therefore also matching number, reduced by one. However, despite the result of Aharoni [Aha2001], this conjecture remained open even for $r=3$ until very recently. As it turns out, with good reason: Clow, Haxell, and Mohar [CHM2025] found two counterexamples for the case $r=3$, namely the line hypergraphs of the Biggs-Smith graph and F168D [^1] A graph6 string and some properties of the F168D graph can be found at http://atlas.gregas.eu/graphs/65., two cubic graphs of respective order $102$ and $168$. Several natural questions arise:
  1. 1.
    Can one can construct an infinite family of counterexamples? [CHM2025, Question 4.1]
  2. 2.
    Are there counterexamples for any $r\geq 4$[CHM2025, Question 4.2]
  3. 3.
    Are the properties of these two graphs, in particular that they are not Cayley, relevant to this application? [CHM2025, Question 4.4].
We give a positive answer to the first question by showing that the line hypergraphs of the generalized Petersen graphs $\operatorname{GP}(5k+11,2)$ form such an infinite family for any $k\geq 0$. Note that the smallest counterexample in this family is of order $33$ coming from a cubic graph of order $22$, see Figure 1, making it significantly smaller and much easier to analyze than the line hypergraph of the Biggs-Smith graph.
[…]
Theorem 2.
There exists an infinite family of cubic graphs whose line hypergraphs are counterexamples to Lovász’ conjecture for $r=3$.
We also give a first answer to the second question by providing several specific counterexamples for the next case $r=4$, likewise derived as the line hypergraphs of regular graphs. These constructions are significantly more complex to describe and the smallest of them has order $140$ coming from a quartic graph of order $70$, see Figure 2.
Theorem 3.
There exist quartic graphs whose line hypergraphs are counterexamples to Lovász’ conjecture for $r=4$.
We can also give an answer to the third question: while the generalized Petersen graphs we use are not Cayley, we did find a specific example of a cubic Cayley graph on $36$ vertices, see Figure 3, as well as a quartic Cayley graph on $96$ vertices. Furthermore, while the Biggs-Smith graph and F168D are both edge-transitive (and the Biggs-Smith graph also distance-transitive), none of our examples are. The Biggs-Smith graph and F168D are also Hamiltonian while the members $\operatorname{GP}(5k+11,2)$ of our infinite family are not whenever $k\equiv 0\mod 6$.
[…]
Figure 3: A $3$-regular Cayley graph in $\mathbb{Z}_{36}$ whose line hypergraph is a counterexample to Lovász’ conjecture for $r=3$.
Finally, let us note that none of the examples we found contradict Ryser’s conjecture or are even extremal for it, a question that has been relevant in its own right [AbuBarPokSza19, HaxNarSza18]. This is unsurprising for $r=3$, as in this case Haxell, Narins, and Szabó [HaxNarSza18] showed that every extremal graph for Ryser’s conjecture also fulfills Lovász’ conjecture. The fact that this also remains true for our counterexamples for $r=4$ indicates that this phenomenon may hold for any uniformity.
Outline. Section 2 recalls basic notions on $r$-uniform hypergraphs, line hypergraphs, and the key independence‐number lemma we will use throughout. Section 3 introduces the family of generalized Petersen graphs $\operatorname{GP}(n,2)$ and proves that their line hypergraphs form infinitely many counterexamples to Lovász’ conjecture for $r=3$, establishing Theorem 2. Section 4 constructs the first quartic counterexamples, establishing Theorem 3, and explains the computational search used to find them. Finally, Section 5 offers brief concluding remarks and open questions.
2 Preliminaries
The line hypergraph $\mathcal{L}=\mathcal{L}(G)$ of an $r$-regular graph $G$ is defined as the $r$-uniform hypergraph whose vertex set is $E_{G}$ and whose hyperedges correspond to the incidence sets in $G$, that is $V_{\mathcal{L}}=E_{G}$ and $e_{v}=\{e_{1},\dots,e_{r}\}\in E_{\mathcal{L}}$ if and only if $\{e_{1},\dots,e_{r}\}$ is the set of all edges incident to a vertex $v\in V_{G}$. Note that $\nu(\mathcal{L})=\alpha(G)$, where $\alpha(G)$ denotes the independence number of $G$.
The main tool needed to establish our results is the following lemma, which was first implicitly stated for the case $r=3$ in [CHM2025, Lemma 2.2] and gives conditions on $G$ under which the line hypergraph $\mathcal{L}$ is a counterexample to Lovász’ conjecture. We include a short proof for the sake of completeness.
Lemma 4.
Let $G$ be an $r$-regular graph satisfying the following two properties:
  1. (i)
    $G$ is $r$-edge-colorable.
  2. (ii)
    Removing any $r-1$ edges along with their vertices does not decrease the independence number.
Then the line hypergraph $\mathcal{L}$ of $G$ is an $r$-uniform counterexample to Conjecture 1.
Proof.
First note that for any $r$-regular graph $G$
$\nu(\mathcal{L})\leq\alpha(G)\;.$ (1)
To show this, let $S(v)=\{e\in E_{G}\mid v\in e\}$ denote the set of edges incident to a given vertex $v$. By definition of the line hypergraph, any matching $M\subseteq E_{\mathcal{L}}$ in $\mathcal{L}$ can be written as $M=\{S(v)\mid v\in I\}$ for some set of vertices $I\subseteq V_{G}$ in $G$. Since $M$ is a matching those $S(v)$ are pairwise disjoint and $I$ therefore an independent set in $G$.
Assume now that $G$ satisfies properties (i) and (ii). Note that $\mathcal{L}$ is $r$-partite since $G$ is $r$-edge-colorable by (i). Let $S\subseteq V_{\mathcal{L}}=E_{G}$ with $|S|=r-1$ be arbitrary but fixed and consider $G^{\prime}=G-(\bigcup_{e\in S}e)$ as well as $\mathcal{L}^{\prime}=\mathcal{L}(G^{\prime})$ and observe that $\mathcal{L}^{\prime}\subseteq\mathcal{L}-S$.
We claim that
$\nu(\mathcal{L}^{\prime})\geq\alpha(G)\;.$ (2)
To see this let $I\subseteq V_{G}$ be the maximum independent set in $G$ given by (ii) for $S$, that is $I\cap\left(\bigcup_{e\in S}e\right)=\emptyset$. For any pair of distinct vertices $v_{1},v_{2}\in I$ the two sets $S(v_{1})$ and $S(v_{2})$ of incident edges are disjoint and therefore $M=\{S(v)\mid v\in I\}$ forms a matching in $\mathcal{L}$. Furthermore, since $I\cap\left(\bigcup_{e\in S}e\right)=\emptyset$, every $S(v)$, $v\in I$, is present as an hyperedge in $\mathcal{L}(G^{\prime})$ and therefore $|M|=|I|$. Combining (1) and (2), we get
[…]
so $\mathcal{L}$ is a counterexample to Lovász’ conjecture. ∎
3 Proof of Theorem 2
Generalized Petersen graphs were introduced in 1950 by Coxeter [C1950] and later given their named in 1969 by Watkins [W1969].
Definition 5.
Given $n\geq 3$ and $1\leq k\leq\lfloor(n-1)/2\rfloor$, the generalized Petersen graph $\operatorname{GP}(n,k)$ is the connected cubic graph with vertex set $V_{\operatorname{GP}(n,k)}=\{a_{i}\mid i\in\mathbb{Z}_{n}\}\cup\{b_{i}\mid i\in\mathbb{Z}_{n}\}$ and edge set
$E_{\operatorname{GP}(n,k)}=\{a_{i}b_{i}\mid i\in\mathbb{Z}_{n}\}\cup\{a_{i}a_{i+1}\mid i\in\mathbb{Z}_{n}\}\cup\{b_{i}b_{i+k}\mid i\in\mathbb{Z}_{n}\}.$
For a given $i\in\mathbb{Z}_{n}$, we call $R_{i}:=\{a_{i},b_{i}\}$ the $i$-th rung. For $i,\ell\in[n]$, we define $S_{i,\ell}:=\bigcup_{i}^{i+\ell-1}R_{j}$ as the union of $\ell$ consecutive rungs starting at $i$.
In order to prove Theorem 2, it suffices to construct an infinite family of cubic graphs satisfying the hypotheses from 4. We will show that the generalized Petersen graphs $\operatorname{GP}(5k+11,2)$, where $k\geq 0$, form such an infinite family. We will use the following two known structural results about the independence number of $\operatorname{GP}(n,2)$.
Theorem 6.
[FGS2011, Theorem 2.5] If $n\geq 5$, then $\alpha(\operatorname{GP}(n,2))=\lfloor 4n/5\rfloor$. In particular, $\alpha(\operatorname{GP}(5k+11,2))=4k+8$.
Lemma 7.
[FGS2011, Lemma 2.4] $\alpha(\operatorname{GP}(n,2)[S_{i,5}])\leq 4$ for any $1\leq i\leq n$.
The main step is encapsulated in the following theorem, which verifies condition (ii) of 4.
Theorem 8.
Let $n=5k+11$ for $k\geq 0$. Then, for any two edges $uv,wx\in E_{\operatorname{GP}(n,2)}$, $\alpha(\operatorname{GP}(n,2)-\{u,v,w,x\})=\alpha(\operatorname{GP}(n,2))$.
The proof of Theorem 8 will be by induction on $k$. The base cases for $n=11,16,21$ are easily verified computationally.
Lemma 9.
If $n\geq 26$, for any two edges $uv,wx\in E_{\operatorname{GP}(n,2)}$, there exists an index $i\in[n]$ such that $\{u,v,w,x\}\cap S_{i,10}=\emptyset$.
The intuition behind this lemma is simple: every edge of $\operatorname{GP}(n,2)$ is contained in a segment of at most 3 consecutive rungs (in the cyclic order of $[n]$), so after removing 2 edges we are left with at least $n-6$ rungs, which are divided in two consecutive segments (or only one, if the two edges overlapped rungs). Hence, if $n-6\geq 19$, one of these segments will be of length at least 10. We will now provide a more rigorous proof.
Proof.
Given a vertex $z\in\operatorname{GP}(n,2)$, let $r(z)$ denote the “rung of $z$”, that is, the $i\in[n]$ such that $z\in R_{i}$. By symmetry of $\operatorname{GP}(n,2)$ we may assume w.l.o.g. that $r(u)=\min\{r(s)\mid s\in\{u,v,w,x\}\}=1$. Then $r(v)\in\{1,2,3\}$. Furthermore, we may assume w.l.o.g. that $r(w)\leq r(x)$.
1. Case: $r(w)>n/2$. Then $r(x)>n/2$ and $I\cap\{r(s)\mid s\in\{u,v,w,x\}\}=\emptyset$ for $I:=\{4,\cdots,n/2\}$. Hence, we can take $i=4$ as the desired index, as $|I|\geq 10$ for $n\geq 26$.
2. Case: $r(w)\leq n/2$. Then $r(x)\leq n/2+2$ and consequently $I\cap\{r(s)\mid s\in\{u,v,w,x\}\}=\emptyset$ for $I:=\{n/2+3,\cdots,n\}$. Hence, we can take $i=n/2+3$ as the desired index, as $|I|\geq 10$ for $n\geq 26$. ∎
Proof of Theorem 8.
The cases $k=0,1,2$ are verified by computational search. Assuming that the statement is true for a certain $k\geq 2$, we will show that the statement is also true for $k+1$. Let $n=5(k+1)+11$, and choose two arbitrary edges $uv,wx\in E_{\operatorname{GP}(n,2)}$. Since $n\geq 26$, by 9 we can find an $i\in[n]$ such that $\{u,v,w,x\}\cap S_{i,10}=\emptyset$. By shifting the labels of the rungs cyclically, we may assume that $i=n-9$.
Consider the graph $\tilde{G}$ with vertices $V(\tilde{G}):=\bigcup_{i=1}^{n-5}R_{i}$ and the edges induced by $\operatorname{GP}(n,2)$ together with edges $a_{n-5}a_{1}$, $b_{n-5}b_{2}$, and $b_{n-6}b_{1}$. Since $\tilde{G}\cong\operatorname{GP}(5k+11,2)$ and by the induction hypothesis, there exists an independent set $\tilde{I}\subseteq V(\tilde{G})$ such that $\{u,v,w,x\}\cap\tilde{I}=\emptyset$ and $|\tilde{I}|=4k+8$.
By 7, $\tilde{I}$ can only intersect at most 4 of $R_{n-9},\dots,R_{n-5}$. Let $i_{0}\in[n-9,n-5]$ be such that $R_{i_{0}}\cap\tilde{I}=\emptyset$. We will construct an independent set $I\subseteq V(G)$ of size $|I|=4k+12$ in the following way:
  1. (i)
    For all $z\in\tilde{I}$ such that $z\in S_{1,i_{0}}$, add $z$ to $I$.
  2. (ii)
    For all $a_{i}\in\tilde{I}$ such that $i>i_{0}$, add $a_{i+5}$ to $I$.
  3. (iii)
    For all $b_{i}\in\tilde{I}$ such that $i>i_{0}$, add $b_{i+5}$ to $I$.
  4. (iv)
    Add $a_{i_{0}+1}$, $b_{i_{0}+2}$, $b_{i_{0}+3}$ and $a_{i_{0}+4}$ to $I$.
Since $\{u,v,w,x\}\subseteq S_{1,i_{0}}$, $\tilde{I}\cap\{u,v,w,x\}=\emptyset$ and $I\cap S_{1,i_{0}}=\tilde{I}\cap S_{1,i_{0}}$ we have $I\cap\{u,v,w,x\}=\emptyset$. Furthermore, $I$ is an independent set in $G$. Indeed, $I\cap S_{1,i_{0}}$ is a independent set in $G[S_{1,i_{0}}]$, since it only includes vertices chosen in (i). $I\cap S_{i_{0},5}$ is a independent set in $G[S_{i_{0},5}]$, since it only includes vertices chosen in (iv). $I\cap S_{i_{0}+5,(n-i_{0}-5+1)}$ is an independent set in $G[S_{i_{0}+5,(n-i_{0}-5+1)}]$, since it only includes vertices chosen in (ii) and (iii). Therefore the only crucial edges left to check are the nine edges connecting the three segments. Since $R_{i_{0}}\cap I=\emptyset$ and $b_{i_{0}+1}\notin I$ none of the three edges $a_{i_{0}}a_{i_{0}+1},b_{i_{0}}b_{i_{0}+2}$, and $b_{i_{0}-1}b_{i_{0}+1}$ are contained in $I$. Similarly, since $R_{i_{0}+5}\cap I=\emptyset$ and $b_{i_{0}+4}\notin I$ none of the three edges $a_{i_{0}+5}a_{i_{0}+6},b_{i_{0}+5}b_{i_{0}+7}$, and $b_{i_{0}+4}b_{i_{0}+6}$ are contained in $I$. Lastly, by (ii) and (iii) none of the edges $a_{n}a_{1},b_{n-1}b_{1}$, and $b_{n}b_{2}$ is contained in $I$, as otherwise one of the edges $a_{n-5}a_{1}$, $b_{n-6}b_{1}$, and $b_{n-5}b_{2}$ would be contained in $\tilde{I}$ contradicting that $\tilde{I}$ is independent in $\tilde{G}$.
Thus $\alpha(\operatorname{GP}(n,2)-\{u,v,w,x\})\geq|I|=4k+12=\alpha(\operatorname{GP}(n,2))$. Since deleting vertices can only decrease the independence number, we conclude that $\alpha(\operatorname{GP}(n,2)-\{u,v,w,x\})=\alpha(\operatorname{GP}(n,2))$.
Proof of Theorem 2.
We need to verify that, for $k\geq 0$, the line hypergraph of $\operatorname{GP}(5k+11,2)$ is a valid counterexample to 1. From 4, it suffices to show that $\operatorname{GP}(5k+11,2)$ is 3-edge-colorable and that it satisfies condition (ii). The 3-edge-colorability has been known since the 70s [Castagna_Prins_1972] and condition (ii) follows from Theorem 8. ∎
Table 1: Overview of counterexamples for $r=3$. $\mathbf{G}^{(36)}_{3}$ has generators $f_{1}$ of order $4$, $f_{2}$ of order $3$, $f_{3}$ of order $2$, and $f_{4}$ of order $3$.
description properties
Biggs-Smith [CHM2025] order $102$, diameter $7$, girth $9$, edge- and distance-transitive, Hamiltonian
F168D [CHM2025] order $168$, diameter $9$, girth $7$, edge-transitive, Hamiltonian
see Figure 3 order $36$, diameter $5$, girth $7$, Cayley graph in $\mathbf{G}^{(36)}_{39}$, Hamiltonian
$\operatorname{GP}(19,7)$ order $38$, diameter $5$, girth $7$, Hamiltonian
$\operatorname{GP}(20,6)$ order $40$, diameter $6$, girth $7$, Hamiltonian
$\operatorname{GP}(25,9)$ order $50$, diameter $6$, girth $7$, Hamiltonian
$\operatorname{GP}(26,8)$ order $52$, diameter $7$, girth $7$, Hamiltonian
$\operatorname{GP}(31,11)$ order $62$, diameter $7$, girth $7$, Hamiltonian
$\operatorname{GP}(32,10)$ order $64$, diameter $8$, girth $7$, Hamiltonian
$\operatorname{CGP}(2,\mathbf{G}^{(36)}_{3};f_{2}\,f_{3}\,f_{4},f_{1}\,f_{2};f_{2}^{2},f_{2})$ order $72$, diameter $6$, girth $6$, Hamiltonian
$\operatorname{GP}(37,17)$ order $74$, diameter $8$, girth $7$, Hamiltonian
$\operatorname{GP}(38,12)$ order $76$, diameter $9$, girth $7$, Hamiltonian
$\operatorname{GP}(43,15)$ order $86$, diameter $9$, girth $7$, Hamiltonian
$\operatorname{GP}(44,14)$ order $88$, diameter $10$, girth $7$, Hamiltonian
$\operatorname{GP}(5k+11,2)$ order $210(k+2)+$ 2, Hamiltonian iff $n\not\equiv 5\mod 6$ [alspach1983classification, Theorem 1]
4 Additional counterexamples
Let us now describe the additional counterexamples beyond those described in the previous section. The properties of all of the graphs have been verified computationally. To describe them, let us define the following family of $3$- and $4$-regular graphs.
Definition 10.
Given $m\geq 2$, some finite multiplicatively-written group $\mathbf{G}$, elements $k_{i}\in\mathbf{G}\setminus{0}$ with $k_{i}\neq(k_{i})^{-1}$, and $c_{i}\in\mathbb{Z}_{m}$ for $i\in\mathbb{Z}_{m}$, the Cayley-generalized Petersen graph $\operatorname{CGP}(m,\mathbf{G};k_{i};c_{i})$ has vertex set $\mathbb{Z}_{m}\times\mathbf{G}$ and connects vertex $(i,j)$ both to $(i,j\,k_{i})$ and $(i+1,j\,c_{i})$ for any $i\in\mathbb{Z}_{m}$ and $j\in G$.
Using this definition, we can list our results for $r=3$ in Table 1 and those for $r=4$ in Table 2. Note that we denote the $i$-th small group of order $n$ by $\mathbf{G}^{(n)}_{i}$ and write it multiplicatively.
Table 2: Overview of counterexamples for $r=4$. The groups $\mathbb{Z}_{28}$, $\mathbb{Z}_{35}$ and $\mathbb{Z}_{51}$ are written additively. The group $\mathbf{G}^{(42)}_{3}$ is $\operatorname{C}_{7}\times S_{3}$ and has generators $f_{1}$ of order $2$, $f_{2}$ of order $7$, and $f_{3}$ of order $3$. The group $\mathbf{G}^{(42)}_{4}$ is $\operatorname{C}_{3}\times\operatorname{D}_{14}$ and has generators $f_{1}$ of order $2$, $f_{2}$ of order $3$, and $f_{3}$ of order $7$. The group $\mathbf{G}^{(48)}_{29}$ is $\operatorname{GL}(2,3)$ and has generators $f_{1}$ of order $2$, $f_{2}$ of order $3$, $f_{3}$ of order $4$, $f_{4}$ of order $4$, and $f_{5}$ of order $2$. The group $\mathbf{G}^{(48)}_{32}$ is $\operatorname{C}_{2}\times\operatorname{GL}(2,3)$ and has generators $f_{1}$ of order $2$, $f_{2}$ of order $3$, $f_{3}$ of order $4$, $f_{4}$ of order $4$, and $f_{5}$ of order 2.
description properties
$\operatorname{CGP}(2,\mathbb{Z}_{35};5,7;15,0)$, see Figure 2 order $70$, diameter $5$, girth $5$, Hamiltonian
$\operatorname{CGP}(2,\mathbf{G}^{(42)}_{3};f_{2}\,f_{3},f_{2}^{2}\,f_{3};f_{1}\,f_{2},1)$ order $84$, diameter $5$, girth $5$, Hamiltonian
$\operatorname{CGP}(2,\mathbf{G}^{(42)}_{4};f_{2}\,f_{3},f_{2}\,f_{3}^{2};f_{1}\,f_{2},1)$ order $84$, diameter $5$, girth $5$, Hamiltonian
$\operatorname{CGP}(3,\mathbb{Z}_{28};18,19,5;18,0,14)$ order $84$, diameter $6$, girth $5$, Hamiltonian
$\operatorname{CGP}(2,\mathbf{G}^{(48)}_{29};f_{2}\,f_{5},f_{2}\,f_{4};f_{1}\,f_{4},1)$ order $96$, diameter $5$, girth $6$, Cayley graph in $\mathbf{G}^{(96)}_{193}$, Hamiltonian
$\operatorname{CGP}(2,\mathbf{G}^{(48)}_{32};f_{1}f_{2},f_{1}f_{2}f_{4};f_{1}f_{2}f_{2}^{2}\,f_{3})$ order $96$, diameter $5$, girth $6$, Cayley graph in $\mathbf{G}^{(96)}_{200}$, Hamiltonian
$\operatorname{CGP}(2,\mathbb{Z}_{51};3,26;47,0)$ order $102$, diameter $6$, girth $5$, Hamiltonian
Let us also give some additional motivation for 10 as well as a description of how we computationally explored constructions based on it. Clow, Haxell, and Mohar [CHM2025] used nauty [mckay1981practical, mckay2014practical] to verify that no cubic graph on at most $20$ vertices yields a counterexample to Lovász’ conjecture. They additional checked all cubic edge-transitive graphs of order up to $300$ [conder2002trivalent, conder2006census, conder2025edge] and determined that the Bigs-Smith graph and F168D are the only two such graphs satisfying the requirements of 4. We initially verified whether the property of being edge-transitive was relevant and if counterexamples based on Cayley graphs might exist by checking the library of vertex-transitive graphs of order at most $47$ [holt2020census], which yielded the cubic $36$-vertex example given in Figure 3.
The library contains no quartic example. Vertex-transitive graphs of order $48$ have been determined [holt2022transitive] but do not appear easily available. We then checked cubic graphs of order $22$ using nauty, which yielded the cubic $22$-vertex example given in Figure 1. The House of Graphs database [Coolsaet2023HoG2] identified it as $\operatorname{GP}(11,2)$ and searching through generalized Petersen graphs suggested the family described in section 3, but did also yield further examples, see Table 1.
Based on this, a natural place to look for further counterexamples to $r=3$ as well as counterexamples for higher uniformities are therefore alternative and even wider reaching generalizations of Petersen graphs. The most notable of these in the literature are permutation graphs and supergeneralized Petersen graphs.
Definition 11 (Permutation graphs [chartrand1967planar, ringeisen1984cycle]).
For a given graph $G$ and a permutation $\alpha$ of its vertices, the permutation graph $\operatorname{P}(G,\alpha)$ consists of two disjoint copies $G_{1}$ and $G_{2}$ of $G$ where vertex $v\in G_{1}$ is connected to $\alpha(v)\in G_{2}$ by an edge. If $G$ is a cycle on $n$ vertices, then the resulting graph is called a cycle permutation graph and denoted by $\operatorname{CP}(n,\alpha)$.
Definition 12 (Supergeneralized Petersen graphs [saravzin2007generalizing]).
Given integers $m\geq 2$, $n\geq 3$, and $k_{i}\in\mathbb{Z}_{n}\setminus{0}$ with $k_{i}\neq-k_{i}$ for $i\in\mathbb{Z}_{m}$, the supergeneralized Petersen graph $\operatorname{SGP}(m,n;k_{i})$ has vertex set $\mathbb{Z}_{m}\times\mathbb{Z}_{n}$ and connects the vertex $(i,j)$ both to $(i,j+k_{i})$ and to $(i+1,j)$ for any $i\in\mathbb{Z}_{m}$ and $j\in\mathbb{Z}_{n}$.
Remark 13.
Note that for $d$-regular $G$ and any permutation, the permutation graph is $(d+1)$-regular. In particular, cycle permutation graphs are $3$-regular. Supergeneralized Petersen on the other hand are either $3$ or $4$-regular if $m=2$ and $4$-regular for $m\geq 3$. Clearly $\operatorname{GP}(n,k)=\operatorname{SGP}(2,n;1,k)$ for any $n$ and $k$ and $\operatorname{GP}(n,k)=\operatorname{CP}(n,\alpha)$ with $\alpha(x)=kx$ if $k$ and $n$ are co-prime.
Permutation graphs are likely to be a very good source of counterexamples to Lovasz’ conjecture, but proved computationally too expensive to explore without any additional assumptions. Supergeneralized Petersen graphs can be easily explored computationally, but proved too restrictive and yielded no counterexamples not already covered by generalized Petersen graphs. 10 constitutes a middle ground between the two, and we explored it by implementing an exhaustive generation of these constructions for given $m$ and $\mathbf{G}$ in Python, relying on SageMath [SageMath2025] and in particular the small groups library [SmallGrp] in GAP [GAP2024] as well as cliquer [Cliquer] to check the conditions of 4 and bliss [JunttilaKaski:ALENEX2007, JunttilaKaski:TAPAS2011] for graph canonization. We mostly explored $m=2$ with groups of order up to around $50$ as well as $m=3$ with groups of order up to around $30$. [^2] At the time of writing the computations were not fully finished and we can therefore give no guarantee of completeness for the larger constructions. A brief exploratory search for $m=4$ and $m=5$ did not yield anything interesting.
5 Concluding remarks
Clow, Haxell, and Mohar [CHM2025, Conjecture 4.2] already asked if a weaker version of the conjecture holds that still would imply Ryser’s conjecture, i.e, if one can always find $k\,(r-1)$ vertices for some $11\leq k\leq r-$ 1 whose removal decreases the matching number by at least $k$. It is unclear at this point if any of the counterexamples described here contradict this weakened version of Lovász’ conjecture. We verified computationally that it still holds for some $r=3$ counterexamples of smaller order. Going to graphs of higher order or considering the counterexamples for $r=4$ may yield different results, but doing so computationally does not seem feasible and our understanding of these constructions at this point is insufficient to decide if the conjecture holds for them purely theoretically.
Acknowledgements
We thank Patrick Morris for bringing this problem to our attention. This research has been partially supported by the Spanish Ministry of Science, Innovation and Universities grant PID2023-147202NB-I00. Aida Abiad is supported by NWO (Dutch Research Council) through the grants VI.Vidi.213.085 and OCENW.KLEIN.475.
The research of Frederik Garbe has been partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through project no. 428212407. Xavier Povill is supported by the Grant PID2023-147202NB-I00 funded by MICIU/AEI/10.13039/501100011033, as well as a FPI-UPC grant from Universitat Politècnica de Catalunya and Banco Santander.
[…]